### v2

# ”Security for Everyone” in Finite Blocklength IRS-aided Systems With Perfect and Imperfect CSI

Abstract
Provisioning secrecy for all users, given the heterogeneity in their channel conditions, locations, and the unknown location of the attacker/eavesdropper, is challenging and not always feasible. The problem is even more difficult under finite blocklength constraints that are popular in ultra-reliable low-latency communication (URLLC) and massive machine-type communications (mMTC). This work takes the first step to guarantee secrecy for all URLLC/mMTC users in the finite blocklength regime (FBR) where intelligent reflecting surfaces (IRS) are used to enhance legitimate users’ reception and thwart the potential eavesdropper (Eve) from intercepting. To that end, we aim to maximize the minimum secrecy rate (SR) among all users by jointly optimizing the transmitter’s beamforming and IRS’s passive reflective elements (PREs) under the FBR latency constraints. The resulting optimization problem is non-convex and even more complicated under imperfect channel state information (CSI). To tackle it, we linearize the objective function, and decompose the problem into sequential subproblems. When perfect CSI is not available, we use the successive convex approximation (SCA) approach to transform imperfect CSI related semi-infinite constraints into finite linear matrix inequalities (LMI). We prove that our proposed algorithm’s converges to a locally optimal solution with low computational complexity thanks to our closed-form linearization approach. This makes the solution scalable for large IRS deployments. Extensive simulations with practical settings show that our approach can ensure secure communication for all users while satisfying FBR constraints even with only imperfect CSI.
Index Terms:
Finite blocklength regime, ultra reliable and low latency communication, intelligent reflective surfaces (IRS), fairness max-min optimization, and secrecy rate (SR).
I Introduction
Intelligent reflective surface (IRS) has emerged as a promising technology, garnering significant interest due to its ability to enhance/tailor the passive radio environment. IRSs consist of passive reflective elements (PREs) equipped with phase shift controllers, allowing them to intelligently manipulate reflected signals [1]. By optimizing signal reflections, IRSs improve reception by creating favorably reflected multi-path signals at receivers. Thanks to their cost-effective design and convenient deployment typically on the facades of high-rise buildings, IRSs hold immense potential for various applications, particularly in urban areas where line-of-sight channels between transmitters (Tx) and receivers (Rx) frequently face obstructions [2]. Particularly, for the Internet of Things (IoT) devices that are limited computing capability and battery/energy, the IRS has gained paramount attention aiming to enhance both spectral and energy efficiency [3].
I-A Related Works and Motivations
Another potential application of IRSs is to enhance the security/privacy of users by purposely manipulating reflected signals from the Tx to facilitate the signal reception at legitimate users while maximizing the multi-user interference/degrading the signals at potential eavesdroppers. In [4], the authors studied the design of a single user IRS-aided system to maximize the legitimate user’s secrecy rate (SR), which is defined as the difference between the rate of the legitimate channel and that of the channel from the transmitter to a potential eavesdropper. The authors maximize the single-user SR by jointly optimizing the beamformer at the transmitter and the IRS. A multi-input-single-output (MISO) IRS-aided system with the presence of multiple eavesdroppers was considered in [5]. The transmitter introduced artificial noise (AN) as a countermeasure to enhance user security. In [6], the authors studied a cell-free IRS-aided network, where they aimed to enhance the user’s secrecy by maximizing the weighted sum of the users’ secrecy rate (WSSR) by jointly optimizing the beamforming vector and the IRS’s PREs. The alternating optimization (AO) algorithm was used to decouple the beamforming factor and the IRS’s PREs, and the resulting problem was tackled using semi-definite relaxation (SDR) and continuous convex approximation. Shi et al. in [7] investigated the SR in a multi-input multi-output (MIMO) IRS-aided system with the presence of a single eavesdropper. When the channel state information (CSI) from the IRS to the users/receivers is unknown or imperfect, the authors in [8] studied an IRS-aided multi-user system, where they formulated a secrecy sum rate (SSR) maximization problem with the eavesdropper’s channel partially known to the receiver. It is worth mentioning that the SSR problem does not guarantee the secrecy for all users.
In parallel, low latency communication (URLLC) and massive machine type communication (mMTC), both relying on finite blocklength (short packet) regime/communication (FBR), have been envisioned as key technologies to serve critical missions of IoT. Since FBR often requires a stricter design approach than the long blocklength regime/communication (LBR) systems[9], maintaining high-reliability communication is more challenging due to the lower channel coding gain. Moreover, in mMTC and URLLC applications such as intelligent transportation, leakage information could expose the user’s location or identity. As a potential solution to secure URLLC and mMTC, the use of IRS has recently attracted paramount interest. IRS links exhibit low latency that is beneficial for the FBR strict requirements [10]. Zhao et al. studied the information freshness in a single user FBR-IRS-aided system with the presence of an eavesdropper in [10]. The authors derived a closed-form expression for the upper bound secrecy outage probability under statistical CSI. The single user’s secrecy outage is minimized by jointly optimizing the blocklength and the IRS PREs. In [11], the authors investigated the problem of transmission power minimization in an IRS-aided system under the FBR regime with the presence of multiple eavesdroppers, by jointly optimizing the beamforming vector, the AN, and the IRS PREs. In [12], the authors derived the outage probability and block error rate (BLER) from users and eavesdroppers. In [13], the authors studied the SSR maximization problem for an FBR-IRS-aided system by jointly optimizing the beamforming vector, the IRS PREs, and the blocklength. The optimization problem was tackled using Riemannian conjugate gradient (RCG) based algorithm. However, the assumption that Alice/Tx can obtain Eve’s CSI, as stated before, is not practical. Note that most earlier work focused only on single-user systems. The SSR maximization approach in LBR-IRS systems is neither applicable to nor can guarantee the secrecy for all users in FBR-IRS systems due to the additional latency requirements of the FBR.
Motivated by the above, this work takes the first step to guarantee secrecy for all URLLC/mMTC users in FBR-IRS-aided systems. Specifically, we consider a popular case where IRSs are used to enhance/aid the signal reception at legitimate users (from a transmitter) under the presence of a potential eavesdropper while the direct/line-of-sight channel between the transmitter to all users and eavesdropper does not exist (e.g., due to blockages). [^1] In reality, it is impractical to consider the underlying problem if the eavesdropper has a line-of-sight channel with the transmitter while the users do not. In such cases, a friendly jammer should be used to protect users’ privacy [14]. We then maximize the minimum SR among all the users while maintaining the FBR constraints by optimizing the transmit beamforming vector and the IRS’s PREs. We consider both cases with perfect and imperfect CSI from the IRS to the users and the eavesdropper. For the perfect CSI, we assume that the transmitter can obtain the correct CSI of all users and the eavesdropper.This is the case when one of the legitimate users gets compromised and acts as an eavesdropper, trying to eavesdrop on other users, hence its CSI would be available to Alice.
When only partial or erroneous CSI (from the IRS to the users and the eavesdropper) is available, the problem is even more challenging due to the semi-infinite constraints invoked by the imperfect CSI and the non-convexity nature of the dispersion factors imposed by the FBR transmission constraints. To tackle the problem, we first introduce slack variables to deal with the transmit beamforming vectors and the IRS’s PREs coupling within the objective function. Secondly, we leverage the successive convex approximation (SCA) technique [15], and the $\mathcal{S}$-procedure [16] to convert the semi-infinite constraints into linear matrix inequalities (LMI). Then, we propose a penalty convex-concave procedure (PCCP) [17] to tackle the unit modulus constraint (UMC) of the IRS. Extensive simulations with practical settings show that maximizing the minimum SR among all the users can achieve a better chance of ensuring secure communications for all users (subject to the location of the eavesdropper) under the FBR constraints even under imperfect CSI conditions.
I-B Contributions
The main contributions are summarized as follows:
  • We take the first step to guarantee secrecy for all URLLC/mMTC users in the finite blocklength regime (FBR) where IRSs are used to enhance legitimate users’ reception while thwarting the potential Eve from intercepting. To that end, we aim to maximize the minimum SR among all users by jointly optimizing the transmitter’s beamforming and IRS’s passive reflective elements (PREs) while meeting FBR latency constraints. The resulting optimization problem is non-convex and even more complicated under imperfect channel state information (CSI).
  • When the CSI is available, we tackle the above non-convex problem by linearizing its objective function with tractable approximation functions, leading to a computationally efficient algorithm. To tackle UMC, unlike traditional methods, i.e. conventional SDR, we directly optimize the PREs arguments to provide a low-complexity solution that can scale with large IRSs. The resulting solution is proved to converge to a locally optimal solution of the original non-convex problem.
  • When perfect CSI is not available, we use the successive convex approximation (SCA) approach to transform imperfect CSI related semi-infinite constraints into finite linear matrix inequalities (LMI). We prove that our proposed algorithm converges to a locally optimal solution with low computational complexity thanks to our closed-form linearization approach. This makes the solution scalable for large IRS deployments.
  • Eventually, we perform extensive simulations with practical settings. The simulations results show that our approach can ensure secure communication for all users while satisfying FBR constraints even under imperfect CSI. Additionally, the impact of the reflected channel’s imperfect CSI on the system is also evaluated.
The remainder of this paper is organized as follows. The system model is discussed in Sections II. Then, Sections III and IV present the problem and corresponding solutions under perfect and imperfect CSI, respectively. The extensive simulations and discussion are in Section V. Finally, Section VI concludes the paper.
This paper uses the following notation: bold letters denote vectors and matrices. $\boldsymbol{I}_{M}$ denotes an M dimensional identity matrix. Diag($n_{1},\dots,n_{n}$) denotes the diagonal matrix with diagonal entries of $\{n_{1},\dots,n_{n}\}$. The symbols $\Re$ and $\mathbb{C}$ represent the real and the complex field, respectively. ${\cal C}(0,\bar{z})$ denotes the circular Gaussian random variable with zero mean and variance $\bar{z}$. For matrices $\mathbf{C}$ and $\mathbf{D}$, $\langle\mathbf{C},\mathbf{D}\rangle\triangleq\mbox{trace}\left(\mathbf{C}^{H}\mathbf{D}\right)$. For matrix $\mathbf{A}$, $\Re\left\{\mathbf{A}\right\}$ denotes the real part, $\langle\mathbf{A}\rangle\triangleq\mbox{trace}(\mathbf{A})$, the symbol $\|\mathbf{A}\|_{1}$ denotes the 1-norm, $\|\mathbf{A}\|$ denotes the Frobenius norm, $\mathbf{A}^{*}$ denotes the conjugate, $\mathbf{A}^{H}$ denotes the Hermitian (conjugate transpose), $\lambda_{\text{max}}\left(\mathbf{A}\right)$ denotes the maximal eigenvalue, $\angle(\mathbf{A})$ denotes its argument, and $\mathbf{A}\succeq 0$ means positive semi-definite.
II System Model
Figure 1: System model.
[…]
$y_{i}\triangleq\boldsymbol{h}_{i}(\boldsymbol{\theta}){\sum}_{k=1}^{\mathcal{K}}\boldsymbol{w}_{k}s_{k}+n_{i},i\in\{k,e\},$ (1)
where $\boldsymbol{h}_{i}(\boldsymbol{\theta})\in\mathbb{C}^{1\times M}$ is the cascaded channel gain from Alice to $i\in\{k,e\}$, $\boldsymbol{w}_{k}\in\mathbb{C}^{M\times 1}$ is the beamforming vector applied for user-$k$, $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{N})^{T}\in[0,2\pi)^{N}$ denotes the phase shift vector the IRS’s PREs, $n_{i}$ is the zero-mean Additive White Gaussian Noise (AWGN) with power density $\sigma_{i}$ for $i\in\{k,e\}$. The cascade channel gain $\boldsymbol{h}_{i}(\boldsymbol{\theta})$ from Alice to $i\in\{k,e\}$, can be written in terms of the channel gain from Alice to the IRS and the channel gain from the IRS to the user or Eve as follows:
$\boldsymbol{h}_{i}(\boldsymbol{\theta})\triangleq{\boldsymbol{u}}_{i}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}\triangleq{\boldsymbol{u}}_{i}\sum_{n=1}^{N}\exp(j\theta_{n})\Xi_{n}\boldsymbol{G}_{{}_{\text{AR}}},$ (2)
where $\boldsymbol{\Xi}_{n}$ is an $N\times N$ matrix with all zeros except for its $(n,n)$ entry which is 1, and $\boldsymbol{\Phi}=\text{Diag}(e^{j\boldsymbol{\theta}})$.
The CSI from Alice to the IRS can be estimated by calculating the angle of arrival and departure [20]. However, the reflected channel’s CSI from the IRS to the IoT users is much more challenging to obtain due to the passive nature of the IRS and the mobility/varying nature of the users’ environment and location [21]. For Eve, it is also hardly possible to pinpoint its location or its accurate CSI from the IRS. To account for the CSI imperfection, we adopt the bounded CSI model in which the reflected channel from the IRS to the users and Eve can be expressed as [22]:
$\boldsymbol{u}_{i}$ $\triangleq\widehat{\boldsymbol{u}}_{i}+\Delta\boldsymbol{u}_{i},\forall i\in\{k,e\},$ (3a)
$\omega_{i}$$\triangleq\{\|\Delta\boldsymbol{u}_{i}\|_{2}\leq\Omega_{i}\},\forall i\in\{k,e\},$ (3b)
where $\widehat{\boldsymbol{u}}_{i}$ denotes the (imperfect) estimated channel vector, $\Delta\boldsymbol{u}_{i}$ represents the channel estimation error of the corresponding estimation, $\omega_{i}$ is a set for all possible channel estimation errors, and $\Omega_{i}$ is the radii of the uncertainty regions as known to Alice. Hence, (2) can be reformulated as:
$\widehat{\boldsymbol{h}}_{i}(\boldsymbol{\theta})\triangleq\left(\widehat{{\boldsymbol{u}}}_{i}+\Delta\boldsymbol{u}_{i}\right)\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}.$ (4)
In the sequel, we deal with perfect and imperfect CSI from the IRS to the users and Eve. For the former, we assume that the CSI of all the users and Eve can be accurately obtained by Alice [23, 24]. As aforementioned, in this case we assume that Eve acts as a legitimate user in one period/slot but then tries to eavesdrop on other users. Using different well-established channel estimation methods, such as the anchor-assisted channel estimation approach [25], Alice can practically obtain the CSI of all the users. In this case, the reflected channel from the IRS to the users and Eve can be captured by setting $\Delta\boldsymbol{u}_{i}$ to zero in (3a), i.e., $\boldsymbol{u}_{i}=\widehat{\boldsymbol{u}}_{i},i\in\{k,e\}$. In the second case, only partial/imperfect CSI $\widehat{\boldsymbol{u}}_{i}$ is available. In this case, one can vary the magnitude of $\Delta\boldsymbol{u}_{i}$, i.e., the radii $\Omega_{i}$ of the uncertainty region to capture different levels of CSI imperfection.
The corresponding signal-to-interference-plus-noise ratio (SINR) of the received signal at the user-$k$ and Eve under perfect and imperfect IRS to users/Eve CSI can be written as, respectively,
$\gamma_{i}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\frac{|\boldsymbol{h}_{i}(\boldsymbol{\theta})\boldsymbol{w}_{k}|^{2}}{\rho_{i}},\leavevmode\nobreak\ i\in\{k,e\},$ (5)
$\widehat{\gamma}_{i}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\frac{|\widehat{\boldsymbol{h}}_{i}(\boldsymbol{\theta})\boldsymbol{w}_{k}|^{2}}{\widehat{\rho}_{i}},\leavevmode\nobreak\ i\in\{k,e\},$ (6)
where $\rho_{i}\triangleq{\sum}_{j=1,j\neq k}^{\mathcal{K}}|\boldsymbol{h}_{i}(\boldsymbol{\theta})\boldsymbol{w}_{j}|^{2}+\sigma_{i}$, and $\widehat{\rho}_{i}\triangleq{\sum}_{j=1,j\neq k}^{\mathcal{K}}|\widehat{\boldsymbol{h}}_{i}(\boldsymbol{\theta})\boldsymbol{w}_{j}|^{2}+\sigma_{i}$. Under the presence of Eve, the closed-form expressions for the FBR-SR of the user-$k$ under perfect and imperfect IRS to users/Eve CSI are defined as, respectively, [26]:
$\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\left[C_{k}-\xi_{k}\sqrt{V_{k}}-\left(C_{e}+\xi_{e}\sqrt{V_{e}}\right)\right]^{+},$ (7)
$\widehat{\mathcal{S}}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\left[\widehat{C}_{k}-\xi_{k}\sqrt{\widehat{V}_{k}}-\left(\widehat{C}_{e}+\xi_{e}\sqrt{\widehat{V}_{e}}\right)\right]^{+},$ (8)
where $[x]^{+}\triangleq\text{max}[0,x]$, while the user’s data rate, and the eavesdropping rate of Eve decoding $s_{k}$, are respectively given by,
$C_{i}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\ln\left(1+\gamma_{i}(\boldsymbol{w},\boldsymbol{\theta})\right),i\in\{k,e\}.$ (9)
$\widehat{C}_{i}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\ln\left(1+\widehat{\gamma}_{i}(\boldsymbol{w},\boldsymbol{\theta})\right),i\in\{k,e\},$ (10)
$\xi_{i}\triangleq{Q^{-1}(\varsigma_{i})}/{\text{ln}(2)\sqrt{N_{t}}}$, $Q^{-1}(\cdot)$ is the inverse of the Gaussian Q-function, $N_{t}\triangleq\mathcal{B}t_{t}$ is the packet length, $\mathcal{B}$ is the bandwidth, $t_{t}$ is the transmission duration, $\varsigma_{k}$ is the decoding error probability[26], $\varsigma_{e}$ is the information leakage[26], and $V_{k}$ and $V_{e}$ are the dispersion factors defined as [26],
$V_{i}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\frac{2\gamma_{i}(\boldsymbol{w},\boldsymbol{\theta})}{(1+\gamma_{i}(\boldsymbol{w},\boldsymbol{\theta}))}\triangleq 2\left(1-({\rho_{i}}/{\upsilon_{i}})\right),i\in\{k,e\},$ (11)
$\widehat{V}_{i}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\frac{2\widehat{\gamma}_{i}(\boldsymbol{w},\boldsymbol{\theta})}{(1+\widehat{\gamma}_{i}(\boldsymbol{w},\boldsymbol{\theta}))}\triangleq 2\left(1-({\widehat{\rho}_{i}}/{\widehat{\upsilon}_{i}})\right),i\in\{k,e\},$ (12)
where $\upsilon_{i}\triangleq{\sum}_{j=1}^{\mathcal{K}}|\boldsymbol{h}_{i}(\boldsymbol{\theta})\boldsymbol{w}_{j}|^{2}+\sigma_{i}$, and $\widehat{\upsilon}_{i}\triangleq{\sum}_{j=1}^{\mathcal{K}}|\widehat{\boldsymbol{h}}_{i}(\boldsymbol{\theta})\boldsymbol{w}_{j}|^{2}+\sigma_{i},\leavevmode\nobreak\ i\in\{k,e\}$.
Unlike the traditional SR definition in the LBR, the FBR induces more constraints on the SR. Specifically, the reliable transmission in FBR requires that the decoding error probability $\varsigma_{k}$ at user-$k$ is not larger than the maximum decoding error probability in FBR $\varsigma_{\max}$. Additionally, the information leakage constraint imposes that the information leakage $\varsigma_{e}$ does not exceed $\varsigma_{\max}$, the maximum information leakage in FBR [26]. These two conditions distinguish the FBR case from its LBR counterpart. One may notice that when the transmission duration $t_{t}$ approaches infinity, the dispersion factor $V_{i}$ reaches zero, and the FBR definitions (7) and (8) reduce to the traditional LBR-SR case, which can be expressed as [27]:
$\mathcal{S}_{k}^{\mathcal{L}}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq[C_{k}(\boldsymbol{w},\boldsymbol{\theta})-C_{e}(\boldsymbol{w},\boldsymbol{\theta})]^{+},$ (13)
$\widehat{\mathcal{S}}^{\mathcal{L}}_{k}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq[\widehat{C}_{k}(\boldsymbol{w},\boldsymbol{\theta})-\widehat{C}_{e}(\boldsymbol{w},\boldsymbol{\theta})]^{+}.$ (14)
III MAXMIN SR under perfect CSI in FBR systems
In this section, we first address the problem of maximizing the minimum SR among all the users under the perfect CSI assumption to provide secure communications for all the users. The optimization problem can be formally stated as:
$(\mathcal{P}1):$ $\leavevmode\nobreak\ \underset{\boldsymbol{w},\boldsymbol{\theta}}{\max}\leavevmode\nobreak\ \underset{k\in\mathcal{K}}{\min}\leavevmode\nobreak\ \mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w},\boldsymbol{\theta}),$ (15a)
s.t. $\leavevmode\nobreak\ {\sum}_{k=1}^{\mathcal{K}}\|\boldsymbol{w}_{k}\|^{2}\leq P,$ (15b)
$|e^{(j\boldsymbol{\theta})}|=1,$ (15c)
[…]
The optimization problem ($\mathcal{P}1$) is non-convex since the objective function (15a) is not concave and the UMC (15c) is non-convex. To tackle this non-convex problem, one can employ the AO technique [1]. Specifically, at iteration $(\iota)$, the feasible point $(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})$ is generated from ($\mathcal{P}1$) by solving two sub-problems, the first one is to optimize $\boldsymbol{w}$ with a fixed $\boldsymbol{\theta}$:
[…]
then, we optimize $\boldsymbol{\theta}$ with a fixed $\boldsymbol{w}$ by solving the following problem
[…]
However, with AO, solving two subproblems $(\mathcal{P}1.1)$ and $(\mathcal{P}1.2)$ is computationally demanding, especially given the large number of PREs of the IRS. Our proposed linearization method in the sequel uses mathematically tractable approximation functions leading to a computationally efficient algorithm that can be used for a large number of IRS’s PREs.
$C_{e}(\boldsymbol{w},\boldsymbol{\theta}^{(\iota)})$ $\geq q_{1,k_{e}}^{(\iota)}+n_{1,k_{e}}^{(\iota)}+2\Re\sum\limits_{j=1,j\neq k}^{\mathcal{K}}\left\{\left\langle\boldsymbol{m}_{j,k_{e}}^{(\iota)},\boldsymbol{w}_{j}\right\rangle\right\}-\frac{\rho_{e}^{(\iota)}}{1+\rho_{e}^{(\iota)}}\left(\sum\limits_{j=1,j\neq k}^{\mathcal{K}}|\boldsymbol{h}_{e}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{w}_{j}|^{2}\right)-\frac{1}{\upsilon_{e}^{(\iota)}}\sum\limits_{j=1}^{\mathcal{K}}|\boldsymbol{h}_{e}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{w}_{j}|^{2},$ (20)
${C_{k}}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta})$ $\geq{q}_{1,k}^{(\iota+1)}+2\Re\left\{{\sum}_{n=1}^{N}{\boldsymbol{m}}^{(\iota+1)}_{k,k}(n)e^{j\theta_{n}}\right\}+\left(e^{j\boldsymbol{\theta}}\right)^{H}{\boldsymbol{\varphi}}_{1,k}^{(\iota+1)}e^{j\boldsymbol{\theta}},$ (24)
$C_{e}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta})$ $\geq q_{1,k_{e}}^{(\iota+1)}+n_{1,k_{e}}^{(\iota+1)}+2\mathfrak{R}\left\{{\sum}_{n=1}^{N}{\boldsymbol{m}}^{(\iota+1)}_{k_{e}}(n)e^{(j\theta_{n})}\right\}+\left(e^{j\boldsymbol{\theta}}\right)^{H}{\boldsymbol{\varphi}}_{1,k_{e}}^{(\iota+1)}e^{j\boldsymbol{\theta}}+\left(e^{j\boldsymbol{\theta}}\right)^{H}{\boldsymbol{\varphi}}_{k_{e},j}^{(\iota+1)}e^{j\boldsymbol{\theta}}.$ (25)
III-1 Sub-Problem for Optimizing the Beamforming Vectors
We fix $\boldsymbol{\theta}$ given $\boldsymbol{w}^{(\iota)}$ and solve the problem ($\mathcal{P}1$) to obtain $\boldsymbol{w}^{(\iota+1)}$ satisfying ${\mathcal{S}}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota)})\geq{\mathcal{S}}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})$. We start by linearizing the objective function in (15a), which consists of four parts: the user-$k$’s SR $C_{k}(\boldsymbol{w},\boldsymbol{\theta})$, Eve’s negative eavesdropping rate $C_{e}(\boldsymbol{w},\boldsymbol{\theta})$, the user-$k$ dispersion factor $V_{k}(\boldsymbol{w},\boldsymbol{\theta})$, and Eve’s dispersion factor $V_{e}(\boldsymbol{w},\boldsymbol{\theta})$ as shown in (7).
First, we adopt the idea in [28] to convert user-$k$’s SR into a linear form. Specifically, using the inequality (74) in Appendix D, let’s define $\boldsymbol{\Lambda}\triangleq\boldsymbol{h}_{k}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{w}_{k}$, ϝ$\boldsymbol{\digamma}\triangleq\rho_{k}$, $\hat{\Lambda}\triangleq\boldsymbol{h}_{k}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{w}_{k}^{(\iota)}$ and ϝ^$\hat{\digamma}\triangleq\rho_{k}$, hence, the users’ data rate can be written as:
$C_{k}(\boldsymbol{w},\boldsymbol{\theta}^{(\iota)})\geq q_{1,k}^{(\iota)}+2\Re\left\{\left\langle\boldsymbol{m}_{k,k}^{(\iota)},\boldsymbol{w}_{k}\right\rangle\right\}-n_{1,k}^{(\iota)}\upsilon_{{}_{k}}^{(\iota)},$ (18)
where
$q_{1,k}^{(\iota)}$ $\triangleq C_{k}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})-\gamma_{k}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})-\sigma_{k}n_{1,k}^{(\iota)},$
$\boldsymbol{m}^{(\iota)}_{k,k}$ $\triangleq{\boldsymbol{h}_{{}_{k}}^{H}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{h}_{{}_{k}}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{w}^{(\iota)}_{k}}/{\rho_{k}^{(\iota)}},$
$n_{1,k}^{(\iota)}$ $\triangleq{1}/{\rho_{k}^{(\iota)}}-{1}/{\upsilon_{{}_{k}}^{(\iota)}}.$
Second, we linearize Eve’s eavesdropping rate, which can be written as:
$-\ln(1+\gamma_{e}(\boldsymbol{w},\boldsymbol{\theta}^{(\iota)}))$ $\triangleq\overset{a_{1}}{\overbrace{\ln(1+\rho_{e}^{(\iota)})}}-\overset{a_{2}}{\overbrace{\ln(1+\upsilon_{e}^{(\iota)})}}.$ (19)
Here, we adopt the idea in [29], where the term ($a_{1}$) in (19) can be linearized by defining $\boldsymbol{z}\triangleq\rho_{e}$ and substituting it in the inequality (75) in Appendix D. The term ($a_{2}$) in (19) can be linearized by defining $\boldsymbol{\Upsilon}\triangleq\upsilon_{e}$ and substituting it in the inequality (71) in Appendix D. Hence, Eve’s eavesdropping rate can be expressed as in (20), where
$q_{1,k_{e}}^{(\iota)}$ $\triangleq\ln\left(\rho_{e}^{(\iota)}\right)-\rho_{e}^{(\iota)}-\ln\left(\upsilon_{e}^{(\iota)}\right)+1,$
$n_{1,k_{e}}^{(\iota)}$ $\triangleq\left(-\frac{\rho_{e}^{(\iota)}}{1+\rho_{e}^{(\iota)}}-\frac{1}{\upsilon_{e}^{(\iota)}}\right)\sigma_{e},$
$\boldsymbol{m}_{k_{e},j}^{(\iota)}$ $\triangleq\boldsymbol{h}_{e}^{H}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{h}_{e}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{w}_{j}^{(\iota)}.$
Next, we linearize the FBR dispersion factors $V_{i}(\boldsymbol{w},\boldsymbol{\theta}),i\in\{k,e\}$. Using (11), the inequalities (72) and (73) in Appendix D, and by defining $x\triangleq V_{i}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta})$, $\mathbf{A}\triangleq\rho_{i}$, and $B\triangleq\varsigma_{i}$, the dispersion factors can be expressed as:
$\xi_{i}\sqrt{V_{i}(\boldsymbol{w},\boldsymbol{\theta}^{(\iota)})}\leq q_{2,i}^{(\iota)}-2\sum\limits_{j=1,j\neq k}^{\mathcal{K}}\Re\left\{\left\langle\boldsymbol{d}_{j,i}^{(\iota)},\boldsymbol{w}_{j}\right\rangle\right\}-n_{2,i}^{(\iota)}\upsilon^{(\iota)},$ (21)
where
$q_{2,i}^{(\iota)}$ $\triangleq\xi_{i}\left(\frac{\sqrt{V_{i}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})}}{2}+\frac{\left(\upsilon^{(\iota)}_{i}\right)^{2}+\rho^{(\iota)}_{i}\sigma_{i}-2\upsilon^{(\iota)}_{i}\sigma_{i}}{\left(\upsilon^{(\iota)}_{i}\right)^{2}\sqrt{V_{i}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})}}\right),$
$\boldsymbol{d}_{j,i}^{(\iota)}$ $\triangleq\frac{\xi_{i}[\boldsymbol{h}_{i}(\boldsymbol{\theta}^{(\iota)})]^{2}\boldsymbol{w}^{(\iota)}_{j}}{\upsilon^{(\iota)}_{i}\sqrt{V_{i}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})}},i\in\{k,e\},$
$n_{2,i}^{(\iota)}$ $\triangleq\frac{\xi_{i}\rho^{(\iota)}_{i}}{\left(\upsilon^{(\iota)}_{i}\right)^{2}\sqrt{V_{i}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})}},i\in\{k,e\}.$
By substitution (18), (20), and (21) into (7), the FBR-SR can be written as:
$\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta}^{(\iota)})$ $\geq{q}_{k}^{(\iota)}+2\Re\left\{\left\langle{\boldsymbol{m}}_{k}^{(\iota)},\boldsymbol{w}_{k}\right\rangle\right\}-(\boldsymbol{w}_{k})^{H}{\boldsymbol{\psi}}^{(\iota)}_{k}\boldsymbol{w}_{k},$ (22)
where
${q}_{k}^{(\iota)}$ $\triangleq q_{1,k}^{(\iota)}+q_{1,k_{e}}^{(\iota)}+n_{1,k_{e}}^{(\iota)}+q_{2,k}^{(\iota)}+q_{2,k_{e}}^{(\iota)},$
${\boldsymbol{m}}_{k}^{(\iota)}$ $\triangleq{\sum}_{j=1}^{\mathcal{K}}\boldsymbol{m}_{j,k}^{(\iota)}+{\sum}_{j=1,j\neq k}^{\mathcal{K}}\left(\boldsymbol{m}_{k_{e},j}^{(\iota)}-\boldsymbol{d}_{j,k}^{(\iota)}-\boldsymbol{d}_{j,k_{e}}^{(\iota)}\right),$
$\boldsymbol{\psi}^{(\iota)}_{k}$ $\triangleq{\sum}_{j=1}^{\mathcal{K}}n_{j}^{(\iota)}\boldsymbol{h}_{j}^{H}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{h}_{j}(\boldsymbol{\theta}^{(\iota)})+n_{k_{e}}^{(\iota)}\boldsymbol{h}_{e}^{H}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{h}_{e}(\boldsymbol{\theta}^{(\iota)}),$
$n_{k}^{(\iota)}$ $\triangleq n_{1,k}^{(\iota)}+n_{2,k}^{(\iota)},$
$n_{k_{e}}^{(\iota)}$ $\triangleq\frac{\rho_{e}^{(\iota)}}{\left(1+\rho_{e}^{(\iota)}\right)}+\frac{1}{\upsilon_{e}^{(\iota)}}+n_{2,k_{e}}.$
$\xi_{i}\sqrt{V_{i}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta})}$ $\leq q_{2,i}^{(\iota+1)}-2\Re\left\{{\sum}_{n=1}^{N}\boldsymbol{d}_{i}^{(\iota+1)}e^{(j\theta_{n})}\right\}+\left(e^{(j\boldsymbol{\theta})}\right)^{H}\boldsymbol{\varphi}_{2,i}^{(\iota+1)}e^{(j\boldsymbol{\theta})},i\in\{k,e\}.$ (26)
$q_{2,i}^{(\iota+1)}$ $\triangleq\frac{\xi_{i}{\left(\left(\upsilon^{(\iota+1)}_{i}\right)^{2}\sqrt{V_{i}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota)})}+2\left(\upsilon^{(\iota+1)}_{i}\right)^{2}+2\sigma_{i}\rho_{i}^{(\iota+1)}-4\sigma_{i}\left(\upsilon^{(\iota+1)}_{i}\right)\right)}}{{\left(2\left(\upsilon^{(\iota+1)}_{i}\right)^{2}\sqrt{V_{i}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota)})}\right)}},$ (27)
$\boldsymbol{d}_{i,j}^{(\iota+1)}(n)$ $\triangleq{\xi_{i}}\bigg{/}{\left(\upsilon^{(\iota+1)}_{i}\sqrt{V_{i}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota)})}\right)}\left(\boldsymbol{w}^{(\iota+1)}_{j}\right)^{H}\boldsymbol{h}_{i}^{H}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{u}_{i}\Xi_{n}{\boldsymbol{G}}_{{}_{\text{{AR}}}}\boldsymbol{w}^{(\iota+1)}_{j},n\in N,j\in K,j\neq k,$ (28)
$q_{k}^{(\iota+1)}$ $\triangleq{q}_{1,k}^{(\iota+1)}+{q}_{1,k_{e}}^{(\iota+1)}+q_{2,k}^{(\iota+1)}+q_{2,k_{e}}^{(\iota+1)}+n_{1,ke}^{(\iota+1)}-\left(e^{j\boldsymbol{\theta}^{(\iota)}}\right)^{H}\boldsymbol{\varphi}_{k}^{(\iota+1)}e^{j\boldsymbol{\theta}}-2\lambda_{\text{max}}\left(\boldsymbol{\varphi}_{k}^{(\iota+1)}\right)N,$ (30)
${\boldsymbol{m}}^{(\iota+1)}_{k}(n)$ $\triangleq{\boldsymbol{m}}^{(\iota+1)}_{k,k}(n)+{\boldsymbol{m}}^{(\iota+1)}_{k_{e}}(n)+{\boldsymbol{d}}^{(\iota+1)}_{k}(n)+{\boldsymbol{d}}^{(\iota+1)}_{k_{e}}(n)+{\sum}_{m=1}^{N}e^{-j\theta_{m}^{(\iota)}}\boldsymbol{\varphi}_{k}^{(\iota+1)}(m,n)+\lambda_{\text{max}}\left(\boldsymbol{\varphi}_{k}^{(\iota+1)}\right).$ (31)
Finally, by substitution (22) in (16), and introducing an auxiliary variable $\Gamma$ as the lower bound of the FBR-SR, we can recast problem ($\mathcal{P}1.1$) as:
$(\mathcal{P}1.3):$ $\max_{\Gamma,\boldsymbol{w}},$ (23a)
s.t. $\Gamma\leq\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta}^{(\iota)}),\leavevmode\nobreak\ \forall k,$ (23b)
($,$ (23c)
problem ($\mathcal{P}1.3$) is an SDP problem with its linearized objective function in (22). This problem can be solved using standard solvers, e.g., the interior point method or the CVX toolbox [30].
III-2 Sub-Problem for Optimizing the PREs
Likewise, given $\boldsymbol{w}$, we aim to find $\boldsymbol{\theta}^{(\iota+1)}$ such that, $\mathcal{S}_{k}(\boldsymbol{w}^{(\iota+1)}_{k},\boldsymbol{\theta}^{(\iota+1)})\geq\mathcal{S}_{k}(\boldsymbol{w}^{(\iota+1)}_{k},\boldsymbol{\theta}^{(\iota)}).$ Similar to the previous section, the lower bound approximation of the user’s SR can be obtained using the inequality (74) in Appendix D. The user’s data rate can be expressed as in (24), where
${q}_{1,k}^{(\iota+1)}$ $\triangleq C_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota)})-\gamma_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota)})-\sigma_{k}{n}_{1,k}^{(\iota+1)},$
${n}_{1,k}^{(\iota+1)}$ $\triangleq{1}/{\rho_{k}^{(\iota+1)}}-{1}/{\upsilon_{{}_{k}}^{(\iota+1)}},$
${\boldsymbol{m}}^{(\iota+1)}_{k,k}(n)$ $\triangleq{\hat{\boldsymbol{m}}^{(\iota+1)}_{k,k}(n)}/{\rho_{k}^{(\iota+1)}},$
$\hat{\boldsymbol{m}}^{(\iota+1)}_{k,k}(n)$ $\triangleq\left(\boldsymbol{w}^{(\iota+1)}_{k}\right)^{H}\boldsymbol{h}_{{}_{k}}^{H}(\boldsymbol{\theta}^{(\iota)}){\boldsymbol{u}}_{k}\boldsymbol{\Xi}_{n}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{w}^{(\iota+1)}_{k},$
$\boldsymbol{\varphi}_{1,k}^{(\iota+1)}$ $\triangleq-{n}_{1,k}^{(\iota+1)}{\sum}_{j=1}^{\mathcal{K}}\boldsymbol{\varphi}_{k,j}^{(\iota+1)},$
$\boldsymbol{\varphi}_{k,j}^{(\iota+1)}$ $\triangleq\left(\boldsymbol{\aleph}_{k,j}^{(\iota+1)}(n)\right)^{*}\boldsymbol{\aleph}_{k,j}^{(\iota+1)}(m),n\in N,m\in N,$
$\boldsymbol{\aleph}_{k,j}^{(\iota+1)}(n)$ $\triangleq\boldsymbol{u}_{k}\boldsymbol{\Xi}_{n}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{w}^{(\iota+1)}_{j}.$
Similarly, we can express Eve’s eavesdropping rate as in (25), where,
$q_{1,k_{e}}^{(\iota+1)}$ $\triangleq\ln\left(\rho_{e}^{(\iota+1)}\right)-\rho_{e}^{(\iota+1)}-\ln\left(\upsilon_{e}^{(\iota+1)}\right)+1,$
$n_{1,k_{e}}^{(\iota+1)}$ $\triangleq\left(-\frac{\rho_{e}^{(\iota+1)}}{1+\rho_{e}^{(\iota+1)}}-\frac{1}{\upsilon_{e}^{(\iota+1)}}\right)\sigma_{e},$
${\boldsymbol{m}}^{(\iota+1)}_{k_{e}}(n)$ $\triangleq{\sum}_{j=1,j\neq k}^{\mathcal{K}}{\boldsymbol{m}}^{(\iota+1)}_{j,k_{e}}(n),$
${\boldsymbol{m}}^{(\iota+1)}_{j,k_{e}}(n)$ $\triangleq\left(\boldsymbol{w}^{(\iota+1)}_{k}\right)^{H}\boldsymbol{h}_{e}^{H}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{u}_{e}\boldsymbol{\Xi}_{n}{\boldsymbol{G}}_{{}_{\text{{AR}}}}\boldsymbol{w}^{(\iota+1)}_{k},$
$\boldsymbol{\varphi}_{1,k_{e}}^{(\iota+1)}$ $\triangleq\left(\frac{-\rho_{e}^{(\iota+1)}}{(1+\rho_{e}^{(\iota+1)})}-\frac{1}{(\sigma_{e}+\upsilon_{e}^{(\iota+1)})}\right)\left(\sum\limits_{j=1}^{\mathcal{K}}{\boldsymbol{\varphi}}_{k_{e},j}^{(\iota+1)}\right),$
$\boldsymbol{\varphi}_{k_{e},j}^{(\iota+1)}$ $\triangleq\left({\boldsymbol{\aleph}}_{k_{e},j}^{(\iota+1)}(n)\right)^{*}\boldsymbol{\aleph}_{k_{e},j}^{(\iota+1)}(m),$
$\boldsymbol{\aleph}_{k_{e},j}^{(\iota+1)}(n)$ $\triangleq\boldsymbol{u}_{e}\boldsymbol{\Xi}_{n}{\boldsymbol{G}}_{{}_{\text{{AR}}}}\boldsymbol{w}^{(\iota+1)}_{j}.$
Next, we express the user-$k$’s and Eve dispersion factor. For $i\in\{k,e\}$, we can express the dispersion factors as in (26), where $q_{i,k}^{(\iota+1)}$ is expressed as in (27), $\boldsymbol{d}_{i,j}^{(\iota+1)}(n)$ is expressed as in (28), and
$\boldsymbol{d}_{i}^{(\iota+1)}$ $\triangleq{\sum}_{j=1,j\neq k}^{\mathcal{K}}\boldsymbol{d}_{i,j}^{(\iota+1)},$
$\boldsymbol{\varphi}_{2,i}^{(\iota+1)}$ $\triangleq{n}_{2,i}^{(\iota+1)}{\sum}_{j=1}^{\mathcal{K}}\boldsymbol{\varphi}_{i,j}^{(\iota+1)},$
${n}_{2,i}^{(\iota+1)}$ $\triangleq{(\xi_{i}\rho^{(\iota+1)}_{i})}/{\left(\left(\upsilon^{(\iota+1)}_{i}\right)^{2}\sqrt{V_{i}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota)})}\right)}.$
By substituting (24), (25), and (26) in (7), the lower bounding concave approximation of the FBR-SR can be expressed as:
$\mathcal{S}_{k}^{\mathcal{F}}\left(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}\right)$ $\geq q_{k}^{(\iota+1)}+2\sum\limits_{n=1}^{N}\Re\left\{\boldsymbol{m}^{(\iota+1)}_{k}(n)e^{j\theta_{n}}\right\},$ (29)
where $q_{k}^{(\iota+1)}$ is expressed as in (30), ${\boldsymbol{m}}^{(\iota+1)}_{k}(n)$ is expressed as in (31), and
$\boldsymbol{\varphi}_{k}^{(\iota+1)}$ $\triangleq\boldsymbol{\varphi}_{1,k}^{(\iota+1)}+\boldsymbol{\varphi}_{1,k_{e}}^{(\iota+1)}+\boldsymbol{\varphi}_{k_{e},j}^{(\iota+1)}+\boldsymbol{\varphi}_{2,k}^{(\iota+1)}+\boldsymbol{\varphi}_{2,k_{e}}^{(\iota+1)}.$
With (29), we formulate the following optimization problem to obtain $\boldsymbol{\theta}^{(\iota+1)}$,
$(\mathcal{P}1.4):$ $\underset{\boldsymbol{\theta}}{\max}\leavevmode\nobreak\ \underset{k\in\mathcal{K}}{\min}\leavevmode\nobreak\ \mathcal{S}_{k}^{\mathcal{F}}\left(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota)}\right),$ (32a)
[…]
To tackle problem ($\mathcal{P}1.4$) we define,
$\theta_{n}^{(\iota+1),k}\triangleq 2\pi-\angle{\boldsymbol{m}}^{(\iota+1)}_{k}(n),n=1,\dots,N,$ (33)
then we find $\boldsymbol{\theta}^{(\iota+1)}$ that maximize the $\mbox{SR}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta})$ and satisfies the UMC constraint (15c),
$\boldsymbol{\theta}^{(\iota+1)}\triangleq\arg\underset{\boldsymbol{\theta}\in\left\{\boldsymbol{\theta}^{(\iota+1),k},k=1,\dots,\mathcal{K}\right\}}{\max}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}).$ (34)
Algorithm 1 Proposed Iterative Algorithm for Solving Problem $(\mathcal{P}1)$
1:Initialize: $(\boldsymbol{w}^{(1)}=\boldsymbol{w}^{*},\boldsymbol{\theta}^{(1)}=\boldsymbol{\theta}^{*})$, convergence tolerance $\epsilon_{t}>0$, and Set $\iota=1$.
2:Repeat
3:Update $\boldsymbol{w}^{(\iota+1)}$ by (23), and $\boldsymbol{\theta}^{(\iota+1)}$ by (34);
4:if $\frac{|\underset{k\in\mathcal{K}}{\min}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota+1)})-\underset{k\in\mathcal{K}}{\min}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})|}{\underset{k\in\mathcal{K}}{\min}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})}\leq\epsilon_{t}$.
5:Then $\boldsymbol{\theta}^{(\iota)}\leftarrow\boldsymbol{\theta}^{(\iota+1)}$, $\boldsymbol{w}^{(\iota)}\leftarrow\boldsymbol{w}^{(\iota+1)}$ and terminate.
6:Otherwise $\iota\leftarrow\iota+1$ and continue.
7:Output $(\boldsymbol{w}^{*}=\boldsymbol{w}^{(\iota)}_{k},\boldsymbol{\theta}^{*}=\boldsymbol{\theta}^{(\iota)})$.
One can notice that by setting $V_{i}=0$, the above FBR optimization problem $(\mathcal{P}1)$ is reduced to the LBR optimization problem, which can be solved similarly.
$\mathbf{X}_{k}\triangleq\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{w}_{k}\boldsymbol{w}_{k}^{(\iota),H}\boldsymbol{G}_{{}_{\text{AR}}}^{H}\mathbf{\Phi}^{(\iota),H}+\mathbf{\Phi}^{(\iota)}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{w}_{k}^{(\iota)}\boldsymbol{w}_{k}^{H}\boldsymbol{G}_{{}_{\text{AR}}}^{H}\mathbf{\Phi}^{(\iota),H}-\mathbf{\Phi}^{(\iota)}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{w}_{k}^{(\iota)}\boldsymbol{w}_{k}^{(\iota),H}\boldsymbol{G}_{{}_{\text{AR}}}^{H}\mathbf{\Phi}^{(\iota),H},$ (40)
$\Delta\boldsymbol{u}_{k}\mathbf{X}_{k}\Delta\boldsymbol{u}_{k}^{H}+2\Re\{(\widehat{\boldsymbol{u}}_{k}^{H}\mathbf{X}_{k})\Delta\boldsymbol{u}_{k}\}+d_{k}\geq(2^{\varphi_{k}}-1)\beta_{k},\forall\|\Delta\boldsymbol{u}_{k}\|_{2}\leq\xi_{k},\forall k.$ (41)
The procedure to solve problem ($\mathcal{P}1$) is described in Algorithm 1, which converges to a locally optimal solution of ($\mathcal{P}1$) as formally stated in the following theorem.
Theorem 1
The obtained solution by Algorithm 1 is a locally optimal solution for problem $(\mathcal{P}1)$.
Proof: See Appendix A.
III-3 Complexity Analysis
The developed Algorithm 1 is designed to tackle problem $(\mathcal{P}1)$ by decoupling the beamforming vector and the IRS’s PREs within the objective function. The resulting problem $(\mathcal{P}1)$ is an SDP problem that can be solved by the interior point method[30]. The algorithm’s complexity can be estimated by its worst-case runtime and the number of the decision variables [31]. Thus, in Algorithm 1, the computational complexity of obtaining $\boldsymbol{w}$ given $\boldsymbol{\theta}$ is $\mathcal{O}(M^{3})$, and the computational complexity of obtaining $\boldsymbol{\theta}$ given $\boldsymbol{w}$ is $\mathcal{O}(N^{3}(N+1))$.
IV MAXMIN SR under Imperfect CSI with FBR systems
To deal with the imperfect CSI from the IRS to the users and Eve, we adopt the channel modeling in equation (3) where only partial/imperfect CSI $\widehat{\boldsymbol{u}}_{i}$ is available. First, we can cast the minimum SR maximization problem as follows:
$(\mathcal{P}2):$ $\underset{\boldsymbol{w},\boldsymbol{\theta}}{\max}\leavevmode\nobreak\ \underset{k\in\mathcal{K}}{\min}\leavevmode\nobreak\ \mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta}),$ (35a)
s.t. ${\sum}_{k=1}^{\mathcal{K}}\|\boldsymbol{w}_{k}\|^{2}\leq P,$ (35b)
$|e^{(j\boldsymbol{\theta})}|=1,$ (35c)
$\|\Delta\boldsymbol{u}_{i}\|_{2}\leq\xi_{i},i\in\{k,e\}.$ (35d)
Unlike the traditional LBR case under imperfect CSI, problem $(\mathcal{P}2)$ is more challenging. First, this is because the additional FBR constraints that are captured by the non-convex dispersion factor expression (12) are embedded within the objective function. Moreover, due to the imperfect CSI as modeled in (3), the semi-infinite constraint (35d) further complicates the problem.
To solve problem $(\mathcal{P}2)$, we first introduce slack variables to decompose the coupling of the beamforming vector and the IRS’s PREs in the objective function to facilitate the AO method. Specifically, we first substitute (13) into (35a), and then we introduce slack variables $z$ as the FBR-SR’s lower bound, ${\varphi}_{k}$ represents the minimum users’ rate, $\mu_{k_{e}}$ represents the maximum eavesdropping rate of Eve, and $\tilde{\varphi}_{i},i\in\{k,e\}$ represents the maximum dispersion factor for the user-$k$ and Eve, respectively. The problem $(\mathcal{P}2)$ can be recast as:
$(\mathcal{P}2.1):$ $\leavevmode\nobreak\ \underset{\boldsymbol{w},\boldsymbol{\theta}}{\max}\leavevmode\nobreak\ z,$ (36a)
s.t. $z\leq{\varphi}_{k}-\mu_{k_{e}}-\tilde{\varphi}_{k}-\tilde{\varphi}_{k_{e}},\forall k,$ (36b)
${\varphi}_{k}\leq\widehat{C}_{k},\forall\|\Delta\boldsymbol{u}_{k}\|_{2}\leq\xi_{k},\forall k,$ (36c)
$\mu_{k_{e}}\geq\widehat{C}_{e},\forall\|\Delta\boldsymbol{u}_{e}\|_{2}\leq\xi_{e},\forall k,$ (36d)
$\tilde{\varphi}_{k}\geq\xi_{k}\sqrt{\widehat{V}_{k}},\forall\|\Delta\boldsymbol{u}_{k}\|_{2}\leq\xi_{k},\forall k,$ (36e)
$\tilde{\varphi}_{k_{e}}\geq\xi_{e}\sqrt{\widehat{V}_{e}},\forall\|\Delta\boldsymbol{u}_{e}\|_{2}\leq\xi_{e},\forall k,$ (36f)
[…]
One can notice that the constraint (35d) in problem $(\mathcal{P}2)$ has been captured by the constraints (36c), (36d), (36e), and (36f) in problem $(\mathcal{P}2.1)$ (referring to the CSI factor in the channel definition (4)). Next, we employ the SCA technique and the $\mathcal{S}$-procedure to transform the semi-infinite non-convex constraints (36c), (36d), (36e), and (36f) to finite LMIs. Then, we leverage a PCCP algorithm to tackle the UMC (35c) [32].
IV-1 Sub-Problem for Optimizing the Beamforming Vectors
To linearize the semi-infinite inequalities in (36c), we first substitute (10) into (36c), hence, (36c) can be expressed as:
$2^{\varphi_{k}}-1\leq\frac{\left|(\boldsymbol{u}_{k}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k}\right|^{2}}{\left\|(\boldsymbol{u}_{k}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{W}_{-k}\right\|^{2}+\sigma_{k}},\forall k.$ (37)
By treating the interference plus noise signal as an auxiliary function $\boldsymbol{\beta}=[\beta_{1},\dots,\beta_{K}]$, (37) can be expressed as:
$\left|(\boldsymbol{u}_{k}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k}\right|^{2}\geq(2^{\varphi_{k}}-1)\beta_{k},\forall k,$ (38a)
$\left\|(\boldsymbol{u}_{k}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{W}_{-k}\right\|^{2}+\sigma_{k}\leq\beta_{k},\forall k.$ (38b)
To circumvent the non-convex semi-infinite inequalities in (38a), we replace the left-hand side of (38a) with their lower bounds using the following Lemma.
Lemma 1
At iteration $(\iota)$, let $\boldsymbol{w}^{(\iota)}$ and $\boldsymbol{\theta}^{(\iota)}$ be the optimal solution, then at the point $(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})$ we can express the linear lower bound of (38a) as:
$\left|(\boldsymbol{u}_{k}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k}\right|^{2}\triangleq\boldsymbol{u}_{k}\mathbf{X}_{k}\boldsymbol{u}_{k}^{H},$ (39)
where $\mathbf{X}_{k}$, is given in (40).
Proof:  Refer to Appendix B.
Next, by substituting (39) in (38a), and using (3a) and Lemma 1, the inequality in (38a) is reformulated as in (41), where $d_{k}\triangleq\widehat{\boldsymbol{u}}_{k}\mathbf{X}_{k}\widehat{\boldsymbol{u}}_{k}^{H}$.
To address the uncertainty of $\{\Delta\boldsymbol{u}_{k}\}$ in (41), we leverage the $\mathcal{S}$-procedure [16].
Lemma 2
($\mathcal{S}$-procedure) For any Hermitian matrix $\mathbf{U}_{i}\in\mathbb{C}^{L\times L}$, vector $\mathbf{u}_{i}\in\mathbb{C}^{L\times 1}$, and scalar $\text{u}_{i}$, for ${i}=0,\dots,Q$. A quadratic function of a variable $x$ is defined as:
$f_{i}(x)\triangleq x^{H}\mathbf{U}_{i}x+2\Re\{\mathbf{u}_{i}^{H}x\}+{\text{u}}_{i}.$ (42)
The condition $f_{0}(x)\geq 0$ such that $f_{i}(x)\geq 0,\leavevmode\nobreak\ i=1,\dots,Q$, holds, if an only if there exists $\mathfrak{n}_{i}\geq 0,i=0,\dots,Q$, such that,
$\begin{bmatrix}\mathbf{U}_{0}&\mathbf{u}_{0}\\ \mathbf{u}_{0}^{H}&\text{u}_{0}\end{bmatrix}-\sum\limits_{i=0}^{Q}\mathfrak{n}_{i}\begin{bmatrix}\mathbf{U}_{i}&\mathbf{u}_{i}\\ \mathbf{u}_{i}^{H}&\text{u}_{i}\end{bmatrix}\succeq 0.$ (43)
Using Lemma 2, we can transform (41) into its equivalent LMIs as:
$\begin{bmatrix}\varpi_{k}\mathbf{I}_{M}+\mathbf{X}_{k}&(\widehat{\boldsymbol{u}}_{k}\mathbf{X}_{k})^{H}\\ (\widehat{\boldsymbol{u}}_{k}\mathbf{X}_{k})&d_{k}-(2^{\varphi_{k}}-1)\beta_{k}-\eta_{k}\Omega_{k}^{2}\end{bmatrix}\succeq 0,\forall k,$ (44)
where $\boldsymbol{\eta}\triangleq[\eta_{1},\dots,\eta_{K}]^{T}\geq 0$ are slack variables. Even though (38a) has been transformed into an LMI form, it is still non-convex due to the non-convexity nature of the term $2^{\varphi_{k}}\beta_{k}$. For that, we adopt the SCA technique [15] to convert the non-convex constraint (38a) to a convex approximation expression. Specifically, performing the first order Taylor approximation, the term $2^{\varphi_{k}}\beta_{k}$ can be upper bounded by:
$(\beta_{k}(2^{\varphi_{k}}))^{ub}\triangleq((\varphi_{k}-\varphi_{k}^{(\iota)})(\beta_{k}^{(\iota)})\ln 2+\beta_{k})2^{\varphi_{k}^{(\iota)}},$ (45)
where $\varphi_{k}^{(\iota)},\beta_{k}^{(\iota)}$ are the value of the variables $\varphi_{k},\beta_{k}$ at iteration $(\iota)$ in the SCA-based algorithm, respectively. Lastly, substituting (45) in (44), the LMIs in (44) can be expressed as:
$\begin{bmatrix}\varpi_{k}\mathbf{I}_{M}+\mathbf{X}_{k}&(\widehat{\boldsymbol{u}}_{k}\mathbf{X}_{k})^{H}\\ (\widehat{\boldsymbol{u}}_{k}\mathbf{X}_{k})&d_{k}-(\beta_{k}(2^{\varphi_{k}}))^{ub}+\beta_{k}-\varpi_{k}\Omega_{k}^{2}\end{bmatrix}\succeq 0,\forall k.$ (46)
Next, we tackle (38b) using Schur’s complement [33].
Lemma 3
(Schur’s complement) For given matrices $\mathbf{U}\succeq 0$, $\mathbf{Y}$, and $\mathbf{Z}$, let a Hermitian matrix $\mathbf{X}$ be defined as:
$\mathbf{X}\triangleq\begin{bmatrix}\mathbf{Z}&\mathbf{Y}^{H}\\ \mathbf{Y}&\mathbf{U}\end{bmatrix}.$ (47)
Then, $\mathbf{X}\succeq 0$ if and only if $\Delta{\mathbf{U}}\succeq 0$, where $\Delta{\mathbf{U}}$ is the Schur’s complement define as $\Delta{\mathbf{U}}\triangleq\mathbf{Z}-\mathbf{Y}^{H}\mathbf{U}^{-1}\mathbf{Y}$.
Using Lemma (3), we can equivalently recast (38b) as:
$\begin{bmatrix}\beta_{k}&\mathbf{t}_{k}^{H}\\ \mathbf{t}_{k}^{H}&\mathbf{I}_{K-1}\end{bmatrix}\succeq 0,\forall\|\Delta\boldsymbol{u}_{k}\|_{2}\leq\Omega_{k},\forall k,$ (48)
where $\mathbf{t}_{k}\triangleq\left((\widehat{\boldsymbol{u}}_{k}^{H}\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\mathbf{W}_{-k}\right)^{H}$, and $\mathbf{W}_{-k}\triangleq[\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{k-1},\boldsymbol{w}_{k+1},\dots,\boldsymbol{w}_{\mathcal{K}}]\in\mathbb{C}^{M\times\mathcal{K}-1}$.
Next, we use the Nemirovski’s Lemma [34] to further handle (48).
Lemma 4
(Nemirovski’s Lemma) For any Hermitian matrix $\mathbf{A}$, matrices $\mathbf{B}$, $\mathbf{C}$, and $\mathbf{X}$, and scalar $t$, the following LMI holds,
$\mathbf{A}\succeq\mathbf{B}^{H}\mathbf{X}\mathbf{C}+\mathbf{C^{H}\mathbf{X}^{H}}\mathbf{B},\leavevmode\nobreak\ \text{for}\|\mathbf{X}\|\leq t,$ (49)
if and only if
$\begin{bmatrix}\mathbf{A}-a\mathbf{C}^{H}\mathbf{C}&-t\mathbf{B}^{H}\\ -t\mathbf{B}&a\mathbf{I}\end{bmatrix}\succeq 0,$ (50)
where $a$ is a non-negative real number.
Using Lemma (4) and introducing the slack variable $\boldsymbol{\varkappa}\triangleq[\varkappa_{1},\dots,\varkappa_{k}]\geq 0$, (48) is recast as:
$\begin{bmatrix}B_{11,k}&\widehat{t}^{H}_{k}&\mathbf{0}_{1\times M}\\ \widehat{t}_{k}&\mathbf{I}_{K-1}&\Omega_{k}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{W}_{-k})^{H}\\ \mathbf{0}_{M\times 1}&\Omega_{k}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{W}_{-k})&\varkappa_{k}\mathbf{I}_{M}\end{bmatrix}\succeq 0,\forall k,$ (51)
where $B_{11,k}\triangleq\beta_{k}-\sigma_{k}-\varkappa_{k}$, and $\widehat{t}_{k}\triangleq\left((\widehat{\boldsymbol{u}}_{k}\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\mathbf{W}_{-k}\right)^{H}$.
Next, we tackle (36d) by firstly substituting (10) in (36d), and then treat the interference plus noise signal in (36d) as an auxiliary function $\boldsymbol{\beta}_{e_{k}}\triangleq[\beta_{e_{1}},\dots,\beta_{e_{k}}]$. We hence can recast (36d) as
$\left|(\boldsymbol{u}_{e}\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k}\right|^{2}\leq(2^{\mu_{k_{e}}}-1)\beta_{ke},\forall k,$ (52a)
$\left\|(\boldsymbol{u}_{e}\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{W}_{-k}\right\|^{2}+\sigma_{e}\geq\beta_{ke},\forall k.$ (52b)
To tackle the uncertainties of $\{\Delta\boldsymbol{u}_{e}\}$ in the constraints (52a) and (52b), we use a similar approach as in (38b). Hence, the equivalent LMIs of (52a) and (52b) are, respectively, given by,
$\begin{bmatrix}C_{e,k}&\widehat{c}_{ke}^{H}&\mathbf{0}_{1\times M}\\ \widehat{c}_{ke}&1&\Omega_{e}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{w}_{k})^{H}\\ \mathbf{0}_{M\times 1}&\Omega_{e}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{w}_{k})&\vartheta_{k}\mathbf{I}_{M}\end{bmatrix}\succeq 0,\forall k,$ (53)
$\begin{bmatrix}C_{11,k}&-\widehat{c}_{ke}^{H}&\mathbf{0}_{1\times M}\\ -\widehat{c}_{ke}&\mathbf{I}_{K-1}&-\Omega_{e}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{W}_{-k})^{H}\\ \mathbf{0}_{M\times 1}&-\Omega_{e}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{W}_{-k})&\varrho_{k}\mathbf{I}_{M}\end{bmatrix}\succeq 0,\forall k,$ (54)
where $\boldsymbol{\vartheta}\triangleq[\vartheta_{1},\dots,\vartheta_{K}]\geq 0$ and $\boldsymbol{\varrho}\triangleq[\varrho_{1},\dots,\varrho_{K}]\geq 0$ are a slack variables,
$C_{e,k}$ $\triangleq(\beta_{ke}(2^{\mu_{k_{e}}}))^{ub}-\beta_{ke}-\vartheta_{k},$
$(\beta_{ke}(2^{\mu_{k_{e}}}))^{ub}$ $\triangleq((\mu_{k_{e}}-\mu_{k_{e}}^{(\iota)})(\beta_{ke}^{(\iota)})\ln 2+\beta_{ke})2^{\mu_{k_{e}}^{(\iota)}},$
$\widehat{c}_{ke}$ $\triangleq((\widehat{\boldsymbol{u}}_{e}^{H}\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\mathbf{W}_{-k})^{H},$
$C_{11,k}$ $\triangleq\beta_{ke}-\sigma_{e}-\varrho_{k}.$
Next, we show the steps to derive the LMI of equations (36e) and (36f). Using (12), the inequalities (72) and (73) in Appendix D, and by defining $x\triangleq V_{i}$, $\mathbf{A}\triangleq\rho_{i}$, and $B\triangleq\varsigma_{i}$, the dispersion factors for $i\in\{k,e\}$ can be expressed as:
Algorithm 2 PCCP Algorithm to solve problem $(\mathcal{P}2.3)$
1:Initialize: $(\boldsymbol{w}^{(1)},\boldsymbol{\theta}^{(1)})$, the maximum penalty factor $o_{\text{max}}$, the scaling factor $\nu\geq 1$, convergence tolerance $\epsilon_{t_{1}}$ and $\epsilon_{t_{2}}$, maximum number of iterations $\iota_{\text{max}}$, and Set $\iota=1$.
2:Repeat
3:For $\iota\leq\iota_{\text{max}}$
4:Solve problem $(\mathcal{P}2.3)$ to obtain $\boldsymbol{\theta}^{(\iota+1)}$;
5:end if $\left\|e^{(j\boldsymbol{\theta}^{(\iota)})}-e^{(j\boldsymbol{\theta}^{(\iota-1)})}\right\|_{1}\leq\epsilon_{t_{1}}$ and $Q\leq\epsilon_{t_{2}}$.
6:else
7:Obtain $o^{(\iota+1)}=\text{min}\{\nu o^{(\iota)},o_{\text{max}}\}$;
8:Update $\iota=\iota+1$;
9:else
10:define a new $\boldsymbol{\theta}^{(1)}$, set $\nu>1$ and $\iota=0$.
11:Output the feasible solution $\boldsymbol{\theta}^{*}=\boldsymbol{\theta}^{(\iota)}$.
$\xi_{i}\sqrt{V_{i}}$ $\leq\frac{\xi_{i}\sqrt{V_{i}^{(\iota)}}}{2}+\frac{\xi_{i}|(\boldsymbol{u}_{i}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k}|^{2}}{\sqrt{V_{i}^{(\iota)}}\left(\|(\boldsymbol{u}_{i}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{W}_{k}\|_{2}^{2}+\sigma_{i}\right)},$ (55)
where $\mathbf{W}_{k}\triangleq[\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{\mathcal{K}}]\in\mathbb{C}^{M\times\mathcal{K}}$. Hence, we can express (36e) and (36f) as,
$\frac{(\tilde{\varphi}_{i}-\mathfrak{L})}{\mathfrak{Y}}\geq\frac{|(\boldsymbol{u}_{i}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k}|^{2}}{\|(\boldsymbol{u}_{i}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{W}_{k}\|_{2}^{2}+\sigma_{i}},\leavevmode\nobreak\ i\in\{k,e\},$ (56)
where $\mathfrak{L}\triangleq\frac{\xi_{i}\sqrt{V_{i}^{(\iota)}}}{2}$ and $\mathfrak{Y}\triangleq\dfrac{\xi_{i}}{\sqrt{V_{i}^{(\iota)}}}$.
Similar to (37), by treating the interference plus noise signal as an auxiliary function $\boldsymbol{\zeta}\triangleq[\zeta_{1},\dots,\zeta_{K}]$, (56) can be expressed as:
$\left|(\boldsymbol{u}_{i}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k}\right|^{2}\leq\left(\frac{(\tilde{\varphi}_{i}-\mathfrak{L})}{\mathfrak{Y}}\right)\zeta_{i},\forall k,i\in\{k,e\},$ (57a)
$\left\|(\boldsymbol{u}_{i}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{W}_{k}\right\|^{2}+\sigma_{i}\geq\zeta_{k},\forall k,i\in\{k,e\}.$ (57b)
Similar to (54), we can express (57b) by its equivalent LMI as:
$\begin{bmatrix}\widetilde{C}_{11,i}&-\tilde{c}^{H}_{i}&\mathbf{0}_{1\times M}\\ -\tilde{c}_{i}&\mathbf{I}_{K}&-\Omega_{i}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{W}_{k})^{H}\\ \mathbf{0}_{M\times 1}&-\Omega_{i}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{W}_{k})&\mathfrak{u}_{i}\mathbf{I}_{M}\end{bmatrix}\succeq 0,j\neq k.$ (58)
where $\widetilde{C}_{11,i}\triangleq\zeta_{k}-\sigma_{k}-\mathfrak{u}_{i}$, $\mathfrak{u}\triangleq[\mathfrak{u}_{1},\mathfrak{u}_{2},\dots,\mathfrak{u}_{i}]\geq 0$ is a slack variable, and $\tilde{c}_{i}\triangleq((\widehat{\boldsymbol{u}}_{i}^{H}\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}))\mathbf{W}_{k})^{H}$.
Next, we can express (57a) as the following LMI:
$\begin{bmatrix}C_{k_{f}}&\tilde{c}^{H}_{ke}&\mathbf{0}_{1\times M}\\ \tilde{d}_{i}&1&\Omega_{e}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k})^{H}\\ \mathbf{0}_{M\times 1}&\Omega_{e}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k})&\tilde{\mathfrak{u}}_{i}\mathbf{I}_{M}\end{bmatrix}\succeq 0,j\neq k.$ (59)
where $C_{k_{f}}\triangleq\frac{\tilde{\varphi}_{k}\zeta_{k}}{\mathfrak{Y}}-\frac{\mathfrak{L}\zeta_{k}}{\mathfrak{Y}}-\tilde{\mathfrak{u}}_{i}$, $\tilde{\boldsymbol{\mathfrak{u}}}\triangleq[\tilde{\mathfrak{u}}_{1},\dots,\tilde{\mathfrak{u}}_{K}]\geq 0$ is a slack variable, and $\tilde{d}_{i}\triangleq((\widehat{\boldsymbol{u}}_{e}\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k})^{H}$.
However, (59) is still non-convex due to the non-convex nature of the term $\tilde{\varphi}\zeta_{k}$. Similar to (44), we employ the SCA technique to tackle the non-convex expression as:
$\left(\tilde{\varphi}\zeta_{i}\right)^{ub}\triangleq\tilde{\varphi}^{(\iota)}\zeta_{i}+(\tilde{\varphi}-\tilde{\varphi}^{(\iota)})\zeta_{i}^{(\iota)}$ (60)
Algorithm 3 Proposed AO Algorithm for Solving Problem $(\mathcal{P}2)$
1:Initialize: $(\boldsymbol{w}^{(1)}=\boldsymbol{w}^{*},\boldsymbol{\theta}^{(1)}=\boldsymbol{\theta}^{*})$, $\iota=1$, and convergence tolerance $\epsilon_{t}>0$.
2:Repeat
3:Update $\boldsymbol{w}^{(\iota+1)}$ by solving $(\mathcal{P}2.2)$, and $\boldsymbol{\theta}^{(\iota+1)}$ by solving problem $(\mathcal{P}2.3)$;
4:if $\frac{|\underset{k\in\mathcal{K}}{\min}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota+1)})-\underset{k\in\mathcal{K}}{\min}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})|}{\underset{k\in\mathcal{K}}{\min}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})}\leq\epsilon_{t}$.
5:Then $\boldsymbol{\theta}^{(\iota)}\leftarrow\boldsymbol{\theta}^{(\iota+1)}$, $\boldsymbol{w}^{(\iota)}\leftarrow\boldsymbol{w}^{(\iota+1)}$ and terminate.
6:Otherwise $\iota\leftarrow\iota+1$ and continue.
7:Output $(\boldsymbol{w}^{*}=\boldsymbol{w}^{(\iota)}_{k},\boldsymbol{\theta}^{*}=\boldsymbol{\theta}^{(\iota)})$.
where $\tilde{\varphi}^{(\iota)}$, $\zeta_{i}^{(\iota)}$ are the value of the variables $\tilde{\varphi}_{i}$, $\zeta_{i}$ at iteration $(\iota)$ in the SCA-based algorithm, respectively. Be substituting (60) in (59), we can express (59) as:
$\begin{bmatrix}\widetilde{C}_{k_{f}}&\tilde{c}^{H}_{ke}&\mathbf{0}_{1\times M}\\ \tilde{d}_{i}&1&\Omega_{e}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k})^{H}\\ \mathbf{0}_{M\times 1}&\Omega_{e}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k})&\tilde{\mathfrak{u}}_{i}\mathbf{I}_{M}\end{bmatrix}\succeq 0,j\neq k,$ (61)
where $\widetilde{C}_{k_{f}}\triangleq\frac{\left(\tilde{\varphi}\zeta_{k}\right)^{ub}}{\mathfrak{Y}}-\frac{\mathfrak{L}\zeta_{k}}{\mathfrak{Y}}-\tilde{\mathfrak{u}}_{i}$.
Eventually, reformulating problem $(\mathcal{P}2.1)$ with (46), (51), (53), (54), (58), and (61) yields the following optimization problem:
$(\mathcal{P}2.2):$ $\leavevmode\nobreak\ \underset{\boldsymbol{w},\boldsymbol{\theta}}{\max}\leavevmode\nobreak\ z$ (62a)
s.t. $z\leq{\varphi}_{k}-\mu_{k_{e}}-\tilde{\varphi}_{k}-\tilde{\varphi}_{k_{e}},\forall k,$ (62b)
[…]
$\boldsymbol{\varpi}\geq 0,\boldsymbol{\varkappa}\geq 0,\boldsymbol{\vartheta}\geq 0,\boldsymbol{\varrho}\geq 0,\mathfrak{u}\geq 0,\tilde{\mathfrak{u}}\geq 0,$ (62d)
[…]
At this point, all nonlinear constraints (36c), (36d), (36e), and (36f) have been linearized as LMIs (46), (51), (53), (54), (58), and (61); thus, problem ($\mathcal{P}2.2$) is an SDP problem that can be efficiently solved using standard solvers, e.g., the interior point method or the CVX toolbox [30].
IV-2 Sub-Problem for Optimizing the PREs
[…]
To apply the PCCP method, we introduce an auxiliary variable set $\mathbf{b}\triangleq\{b_{n}|n\in N\}$ satisfying $b_{n}\triangleq|e^{(j\theta_{n})}|^{*}|e^{(j\theta_{n})}|$. Then, (35c) can be expressed as $b_{n}\leq|e^{(j\theta_{n})}|^{*}|e^{(j\theta_{n})}|\leq b_{n}$. The non-convex part $b_{n}\leq|e^{(j\theta_{n})}|^{*}|e^{(j\theta_{n})}|$ can be approximated by $b_{n}\leq 2\Re\{|e^{(j\theta_{n})}|^{*}|e^{(j\theta^{(\iota)}_{n})}|-|e^{(j\theta^{(\iota)}_{n})}|^{*}|e^{(j\theta^{(\iota)}_{n})}|\}$. Following the PCCP framework, we penalize the objective function (36a), hence, problem $(\mathcal{P}2.1)$ can be recast as:
$(\mathcal{P}2.3):$ $\underset{\boldsymbol{\theta}}{\max}\leavevmode\nobreak\ z-oQ,$ (63a)
[…]
$|e^{(j\theta_{n})}|^{*}|e^{(j\theta_{n})}|\leq b_{n}+c_{n},$ (63c)
$b_{n}-\hat{c}_{n}\leq 2\Re\left\{|e^{(j\theta_{n})}|^{*}|e^{(j\theta^{(\iota)}_{n})}|\right\}$
$-|e^{(j\theta^{(\iota)}_{n})}|^{*}|e^{(j\theta^{(\iota)}_{n})}|,$ (63d)
$b_{n}\geq 0,\forall n\in N,$ (63e)
where $Q\triangleq{\sum}_{n=1}^{N}c_{n}+\hat{c}_{n}$ is the penalty term, $\mathbf{c}\triangleq\{c_{n},\hat{c}_{n}\}$ is the slack variable imposed over the modulus constraint (35c), and $o$ is the regularization factor. The regularization factor is imposed to control the UMC constraint (35c) by scaling the penalty term $Q$.
Problem $(\mathcal{P}2.3)$ is an SDP problem that can be solved by the CVX toolbox. Unlike the conventional SDR method to tackle UMC, the PCCP method, summarized in Algorithm 2, is guaranteed to find a feasible point for problem $(\mathcal{P}2.3)$ [35]. The following points are helpful for the numerical implementation of the Algorithm 2:
  1. (a)
    We invoke $o_{\text{max}}$ to avoid numerical complications if $o$ grows too large[36];
  2. (b)
    The stopping criteria $Q\leq\epsilon_{t_{2}}$ guarantees the UMC (35c) when $\epsilon_{t_{2}}$ is small [37];
  3. (c)
    The convergence of Algorithm 2 is controlled by $\left\|e^{(j\boldsymbol{\theta}^{(\iota+1)})}-e^{(j\boldsymbol{\theta}^{(\iota)})}\right\|_{1}\leq\epsilon_{t_{1}}$.
Finally, the pseudo code of the AO algorithm to solve problem ($\mathcal{P}2$) is shown in Algorithm 3. This algorithm converges to a locally optimal solution of ($\mathcal{P}2$), which can be proved in the following theorem.
Theorem 2
Algorithm 3 generates a locally optimal solution for problem $(\mathcal{P}2)$.
Proof: See Appendix C.
IV-3 Complexity Analysis
[…]
$\mathcal{O}_{\boldsymbol{w}}$ $\triangleq\mathcal{O}\left\{\sqrt{{b_{1}}+2}({c})({c}^{2}+{c}{b_{2}}+{b_{3}}+({c}+1)^{2})\right\},$ (64)
$\mathcal{O}_{\boldsymbol{\Phi}}$ $\triangleq\mathcal{O}\left\{\sqrt{{b_{1}}+4N}(2N)(4N^{2}+2N{b_{2}}+{b_{3}}+4NM)\right\},$ (65)
where
${b_{1}}$ $\triangleq{\sum}_{k}^{\mathcal{K}}({a_{1}}+{a_{2}})+2k({a_{2}}+{a_{3}})+(k-2)({a_{1}}+{a_{3}}),$
${b_{2}}$ $\triangleq{\sum}_{k}^{\mathcal{K}}({a_{1}}^{2}+{a_{2}}^{2})+2k({a_{2}}^{2}+{a_{3}}^{2})+(k-2)({a_{1}}^{2}+{a_{3}}^{2}),$
${b_{3}}$ $\triangleq{\sum}_{k}^{\mathcal{K}}({a_{1}}^{3}+{a_{2}}^{3})+2k({a_{2}}^{3}+{a_{3}}^{3})+(k-2)({a_{1}}^{3}+{a_{3}}^{3}).$
V Numerical Results
TABLE I: Numerical Parameters
Parameter Numerical Value
Number of IRS elements ($N$) 16
Noise power density ($\sigma_{i}$) -174 dBm/Hz
Alice Transmission power ($P$) 20 dBm
Antennas’ Gain $(G_{\text{\text{A}}})$ and $(G_{\text{IRS}})$ 5 dBi
The convergence tolerance $\epsilon_{t}$ $10^{-3}$
Sim. initial settings $\beta_{k}^{(\iota)}=1$, $\beta_{ke}^{(\iota)}=1$, $o^{(1)}=10$, $o_{\text{max}}=30$, $\xi_{k}=\xi_{e}=0.02$.
In this section, we perform an extensive simulation to evaluate the performance of the proposed approach. The results are obtained using the MATLAB and CVX toolboxes. In this setup, Alice is located at $(15,0,15)$, the IRS is located at $(0,25,40)$, and the users are randomly distributed to the right of Alice over a $(60m\times 60m)$ area. The Eve (e.g., a compromised legitimate user) is randomly located in $(100m\times 100m)$ outside the users’ area. Note that if Eve is too close to one of the users, it is mainly impossible to guarantee the positive SR for all the users [38]. In this case, other methods, e.g., encryption or friendly jamming, can provide users’ secrecy [14].
The Alice-to-IRS direct path loss factor is ${\beta_{{}_{\text{{AR}}}}}\triangleq\boldsymbol{G}_{\text{\text{A}}}+\boldsymbol{G}_{\text{IRS}}-35.9-22\log_{10}(d_{{}_{\text{{AR}}}})$ in dB, where $d_{{}_{\text{{AR}}}}$ is the distance between Alice and the IRS in meters, while $\boldsymbol{G}_{\text{\text{A}}}$ is Alice antenna gain, and $\boldsymbol{G}_{\text{IRS}}$ is the IRS elements’ antenna gain [39]. The path loss factor from the IRS to the users and the Eve is ${\beta_{\text{R}i}}\triangleq\boldsymbol{G}_{\text{IRS}}-33.05-30\log_{10}(d_{\text{R}i})$ dB, for $i\in\{k,e\}$, $d_{\text{R}i}$ is the distance between the IRS and the users and Eve in meters. The spatial correlation matrix is $[\boldsymbol{R}_{Ri}]_{l,\bar{l}}\triangleq\exp{(j\pi(l-\bar{l})\sin\hat{\vartheta}\sin\hat{\aleph})}$ for $i\in\{k,e\}$, where $\hat{\aleph}$ is the elevation angle and $\hat{\vartheta}$ is the azimuth angle [19]. The elements of the Alice-to-IRS channel are generated by $[\boldsymbol{G}_{{}_{\text{{AR}}}}]_{a,b}\triangleq\exp({j\pi\left((b-1)\sin\overline{\Theta}_{b}\sin\overline{\vartheta}_{b}+(a-1)\sin(\Theta_{n})\sin(\vartheta_{b})\right)})$, where ${\Theta}_{n}\in(0,2\pi)$, ${\vartheta}_{n}\in(0,2\pi)$, and $\overline{\Theta}_{n}\triangleq\pi-\Theta_{n}$, and $\overline{\vartheta}_{n}\triangleq\pi+\vartheta_{n}$ [19]. The small-scale fading channel gain $\widehat{\boldsymbol{u}}_{i}$ for $i\in\{k,e\}$ is modeled as a Rician fading channel with K-factor$=3$.
We define the CSI error bounds as $\Omega_{i}\triangleq\delta_{i}\|\widehat{\boldsymbol{u}}_{i}\|_{2},\forall k$, where $\delta_{i}\in[0,1),i\in\{k,e\}$ is the relative amount of CSI uncertainty. When $\delta=0$, Alice can obtain the perfect CSI of the IRS to users/Eve reflected channel, the transmission duration is $t_{t}=0.1$ms which is suitable for FBR transmission [40], and the choice of $0.1$ ms end-to-end delay ensures having a quasi-static channel during FBR communication [41], $\xi_{i}=10^{(-5)}$, the bandwidth $B=1$Mhz. Unless stated otherwise, the simulation parameters are defined in Table I. We multiply the results by $\log_{2}(e)$ to convert them to bps/Hz. Lastly, for comparison purposes, we compare our Max-Min algorithm against the SSR maximization algorithms. Due to the page limitation, the steps to maximize SSR has been omitted, however, the SSR maximization can be tackled by setting the objective function in problem ($\mathcal{P}1$) and problem ($\mathcal{P}2$) to $\underset{\boldsymbol{w},\boldsymbol{\theta}}{\max}\sum_{k}^{\mathcal{K}}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta})$.
Figure 2: Convergence rate of the Max-Min algorithm under prefect CSI with $M=10$, $\mathcal{K}=6$, $N=16$.
Fig. 2 illustrates the convergence rate of the Max-Min algorithms under perfect CSI and Imperfect CSI for the FBR and the LBR systems. All algorithms achieve convergence within a few numbers of iterations. It can be noticed that the FBR needs fewer iterations to converge since the local optimal solution obtained in the LBR case is used as the initial feasible point. Such choice highlights the importance of choosing the best initial feasible points in the FBR case.
Fig. 3 plots the users’ SR distribution with $M=10$, $\mathcal{K}=6$, and $N=16$ under the Max-Min algorithms. The Max-Min algorithms achieve almost the same SR for all the users when the IRS’s PREs are optimized. However, when the IRS’s PREs are not practically optimized, the algorithms fail to achieve secrecy fairness among the users. The results demonstrate the importance of properly optimizing the IRS’s PREs. The proposed algorithm for the FBR case achieves secrecy and SR fairness among the users. The results demonstrate the robustness of the proposed algorithms to achieve secrecy for all the users, even under the FBR constraints, namely, the transmission duration and the latency.
Figure 3: Users’ SR with $M=10$, $\mathcal{K}=6$, $N=16$.
Figure 4: Users’ SR with with $M=10,\mathcal{K}=6,N=16$.
Fig. 4 portrays the users’ SR under different values of the IRS to the users/Eve imperfect CSI. The Max-Min algorithm achieves secrecy fairness in the LBR/FBR cases, even when Eve’s reflected channel has higher uncertainty than the users’ reflected channel.
Figure 5: Jain’s Index Vs. the uncertainty level $\delta_{k}=\delta_{e}$ with $M=10,\mathcal{K}=5,N=16$.
[…]
$\ln\left|I_{n}+[\boldsymbol{\Lambda}]^{2}(\boldsymbol{\digamma})^{-1}\right|\geq\ln\left|I_{n}+[\hat{\Lambda}]^{2}(\hat{\digamma})^{-1}\right|-\langle[\hat{\Lambda}]^{2}(\hat{\digamma})^{-1}\rangle+2\Re\{\langle\hat{\Lambda}^{H}(\hat{\digamma})^{-1}\mathbf{A}\rangle\}-\langle(\hat{\digamma})^{-1}-(\hat{\digamma}+[\hat{\Lambda}]^{2})^{-1},[\mathbf{A}]^{2}+\boldsymbol{\digamma}\rangle,$ (74)
$\ln(1+\sum_{i=1}^{l}|{z}_{i}|^{2})\geq\ln(1+\sum_{i=1}^{l}|\bar{z}_{i}|^{2})-\sum_{i=1}^{l}|\bar{z}_{i}|^{2}+\sum_{i=1}^{l}2\Re\{\bar{z}_{i}^{*}{z}_{i}\}-\frac{\sum_{i=1}^{l}|\bar{z}_{i}|^{2}\left(1+\sum_{i=1}^{l}|{z}_{i}|^{2}\right)}{1+\sum_{i=1}^{l}|\bar{z}_{i}|^{2}}.$ (75)
Fig. 6 illustrates the minimum SR among all users against the relative amount of CSI uncertainty $\delta$. The Max-Min algorithm achieves secrecy for all the users. Even under high uncertainty, it achieves almost the same minimum SR as the full certainty case. The FBR case demonstrates results similar to those of the LBR case. As expected, the SSR-LBR/SSR-FBR algorithms achieve a higher minimum user SR since the SSR algorithm favors most of the transmission power towards the user with a better channel while discarding other users.
Figure 6: Minimum SR Vs. the uncertainty level $\delta_{k}=\delta_{e}$ with $M=10,\mathcal{K}=5,N=16$.
Figure 7: Minimum SR Vs. the number of users $\mathcal{K}$ with $M=10,N=16$, $\delta_{k}=\delta_{e}=0.02$.
Fig. 7 depicts the minimum SR against the number of users $\mathcal{K}$. It can be seen that the minimum SR algorithm decreases with $\mathcal{K}$. This is expected as the multi-user interference increases with $\mathcal{K}$. The Max-Min FBR case provides secrecy in all cases; however, as $\mathcal{K}$ increases, the minimum users’ SR is approaching zero, which can be seen when $\mathcal{K}=7$. The results show that obtaining secrecy for all the users in the FBR case is more challenging due to the FBR constraints.
Figure 8: user’s min SR Vs. Transmission Duration $t_{t}$.
Lastly, the minimum SR is illustrated against the transmission duration $t_{t}$ in Fig. 8. The Max-Min algorithm guaratnees secure communication for all the users even at low transmission duration ($10^{-5}$ms). It can be noticed that the achievable FBR’s SR is capped by its LBR counterpart since the transmission duration is considered to be infinity in the latter.
VI Conclusions
In this paper, we proposed a framework to achieve secure communications for all the users under the FBR constraints for IoT setting in URLLC/mMTC applications aided by IRS. Specifically, through linearization and different non-convex optimization techniques, we designed computationally efficient algorithm to maximize the minimum SR among all users under both perfect and imperfect CSI from IRS to the users and the eavesdropper. Extensive simulations results showed that the Max-Min algorithm can provide secure communications for all the users under FBR constraints even with only imperfect CSI. Note that if Eve is too close to one user, it is infeasible to guarantee secrecy for all the users and other solutions like cryptography-based or friendly jamming-based methods should be in place. In the future, one can consider a multi-hop scenario with the joint coding over multiple IRSs, similar to the cooperative MIMO setting [43].
Appendix A Proof of theorem 1
First we prove that the sequence $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota+1)})$ is non decreasing for all $k$, i.e. $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota+1)})\geq\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})$ for all $\iota>0$. To that end, with the aid of the CVX solver, we obtain $\boldsymbol{w}^{(\iota+1)}$ by solving problem ($\mathcal{P}1.3$). CVX solver is guaranteed to find the optimal solution. Thus, we have $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta})>\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta})$ for all $k$. Next, from ($\mathcal{P}2$), we obtain $\boldsymbol{\theta}^{(\iota+1)}$ by solving (33), where we have $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w},\boldsymbol{\theta}^{(\iota+1)})>\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w},\boldsymbol{\theta}^{(\iota)})$ as stated before.
Hence, by combining the solutions obtained by solving problem ($\mathcal{P}1.3$) and problem ($\mathcal{P}1.4$), we obtain:
$\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota+1)})>\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)}),$ (66)
the optimal sequence $\{(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota+1)})\}$ thus converge to a point $\{({\boldsymbol{w}}^{*},{\boldsymbol{\theta}}^{*})\}$ which is the solution obtained from solving ($\mathcal{P}1.3$) and ($\mathcal{P}1.4$).
Next, we prove that the converged point $\hat{X}^{*}\triangleq\{\boldsymbol{w}^{*},\boldsymbol{\theta}^{*}\}$ is a locally optimal solution of problem $(\mathcal{P}1)$. For that, we show that the converged point satisfies the Karush-Kuhn-Tucker (KKT) condition of the problem. The KKT condition for problem $(\mathcal{P}1.3)$ is satisfied at $\boldsymbol{\theta}^{*}$. Let $\boldsymbol{Y}(\boldsymbol{\theta})$ be the objective function of $(\mathcal{P}2)$ and $\boldsymbol{T}({\boldsymbol{X}})=[\boldsymbol{T}_{1}({\boldsymbol{X}}),\boldsymbol{T}_{2}({\boldsymbol{X}}),\dots,\boldsymbol{T}_{I}({\boldsymbol{X}})]$ be the set of constraint of problem $(\mathcal{P}2)$. Then, we can write:
$\nabla_{\boldsymbol{\theta}^{*}}\boldsymbol{Y}({\boldsymbol{X}}^{*})+\boldsymbol{Z}^{T}\nabla_{\boldsymbol{\theta}^{*}}\boldsymbol{T}({\boldsymbol{X}}^{*})=0,$ (67)
$z_{i}\geq 0,t_{i}\boldsymbol{T}({\boldsymbol{X}}^{*})=0,\forall i.$
where $\nabla_{s}$ is the gradient with respect to $s$, and $\boldsymbol{Z}=[z_{1},z_{2},\dots,z_{I}]$ is the optimal Lagrangian variable set. Similarly, the obtained $(\mathcal{P}2)$ solution is locally optimal. Hence, its KKT is satisfied with respect to $\boldsymbol{\text{W}}=\boldsymbol{w}_{k}^{*}$, which is:
$\nabla_{\boldsymbol{\text{W}}^{*}}\boldsymbol{Y}({\boldsymbol{X}}^{*})+Z^{T}\nabla_{\boldsymbol{\text{W}}^{*}}\boldsymbol{T}({\boldsymbol{X}}^{*})=0,$ (68)
$z_{i}\geq 0,t_{i}\boldsymbol{T}({\boldsymbol{X}}^{*})=0,\forall i.$
Combining (67) and (68), we get:
$\nabla_{{\boldsymbol{X}}^{*}}\boldsymbol{Y}({\boldsymbol{X}}^{*})+Z^{T}\nabla_{{\boldsymbol{X}}^{*}}\boldsymbol{T}({\boldsymbol{X}}^{*})=0,$ (69)
$t_{i}\geq 0,t_{i}\boldsymbol{T}({\boldsymbol{X}}^{*})=0,\forall i,$
which is the KKT condition for $(\mathcal{P}1)$ in (15a), i.e., $\hat{\boldsymbol{X}}^{*}$ is the local optimal point for (15a). $\blacksquare$
Appendix B Proof of Lemma 1
Let $q$ be a scalar complex variable, and $q^{(\iota)}$ is the fixed point obtained at iteration $(\iota)$, then the following inequality holds [20]:
$|q|^{2}\geq q^{*,(\iota)}q+q^{*}q^{(\iota)}-q^{*,(\iota)}q^{(\iota)}.$ (70)
By replacing $q$ with $({\boldsymbol{u}}_{k}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k}$ we obtain (39). Thus, this concludes the proof. $\blacksquare$
Appendix C Proof of theorem 2
Similar to the proof of Theorem 1, it can be shown the sequence $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota+1)})$ is non-decreasing for all $k$, i.e. $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota+1)})\geq\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})$ for all $\iota>0$, and the converged point is a locally optimal solution for problems. $\blacksquare$.
Appendix D Inequalities
Here, we adopt the inequality (48) in[28]. Specifically, for any $\mathbf{A}$ and $\boldsymbol{\digamma}$ with $\hat{\Lambda}$ and $\hat{\digamma}$ are fixed point, (74) holds.
Next, we adopt Lemma (2) in [29]. Specifically, for any ${z}_{i}$ with $i=1,\dots,l$ and $\bar{z}_{i}$ is a fixed point, inequality (75) holds.
Lastly, the concave logarithmic function can be written as [29, Eq. 15]:
$-\ln(1+\boldsymbol{\Upsilon})\geq-\ln(1+\bar{\Upsilon})-\frac{1+\boldsymbol{\Upsilon}}{1+\bar{\Upsilon}}+1.$ (71)
The following inequality follows from the concavity of the function $\sqrt{x}$ [2]:
$\sqrt{x}\leq{\sqrt{\bar{x}}}/{2}\left(1+{x}/{\bar{x}}\right),\leavevmode\nobreak\ \forall x>0,\bar{x}>0.$ (72)
Lastly, for a any vector $\mathbf{A}\in\mathbb{C}^{n}$, $\bar{\mathbf{A}}\in\mathbb{C}^{n}$, scalar $B>0,\leavevmode\nobreak\ \bar{B}>0,\leavevmode\nobreak\ \sigma>0$, The following inequality (73) holds [2]:
$\frac{||\mathbf{A}||^{2}}{B+\sigma}\geq\frac{||\bar{\mathbf{A}}||^{2}}{\bar{B}+\sigma}\left(2\frac{\Re\{\bar{\mathbf{A}}^{H}\mathbf{A}\}}{||\bar{\mathbf{A}}||^{2}}-\frac{B+\sigma}{\bar{B}+\sigma}\right).$ (73)